OpenStudy (anonymous):

The population of Preston is 89,000 and is decreasing by 1.8% each year. Write a function that models the population as a function of time t.

OpenStudy (immanuelv):

89000 is the constant here that means it wont change and since the population is decreasing the slope will be negative so f(t) = -.018t +89000

OpenStudy (anonymous):

\[ f(t) = 89000 (1-1.8\%)^t = 89000 (1-0.018)^t =89000 (0.982)^t \]

OpenStudy (anonymous):

Thanks !